OpenStudy (anonymous):

Jeannie is 8 years older than her younger sister April. When added together, their ages equal 198. How old is April?

OpenStudy (whpalmer4):

let \(j\) be Jeannie's age, and \(a\) be April's age. \[j = 8+a\]because Jeannie is 8 years older than April \[a+j = 198\]because their ages add to 198 substitute the previous equation: \[a+(8+a) = 198\] Solve for \(a\) to find April's age

OpenStudy (anonymous):

April is 95